Suds & Cuts is a local pet grooming shop owned by Collin Bark. Collin has prepared the following standard cost card for each dog bath given: Standard Quantity Standard Rate Standard Unit Cost Shampoo 1.3 oz. $ 0.30 per oz. $ 0.39 Water 11 gal. $ 0.04 per gal. 0.44 Direct labor 0.85 hr. $ 9.00 per hr. 7.65 During the month of July, Collin’s employees gave 310 baths. The actual results were 705 ounces of shampoo used (cost of $232.65), 6,100 gallons of water used (cost of $427), and labor costs for 250 hours (cost of $2,600). Required: 1. Calculate Suds & Cuts direct materials variances for both shampoo and water for the month of July. (Round your answers to 2 decimal places. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable.) 2. Calculate Suds & Cuts direct labor variances for the month of July. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable. Round your answers to nearest dollar amount.)
Answer #1

1. Materials price variance = (Standard price - actual price) x Material purchased

Shampoo = {$0.30 - ($232.65/705)} x 705 = $21 Unfavorable

Water = ($0.04 - 427/6100)*6100 = $183 unfavorable

Materials quantity variance = (Standard quantity - actual quantity used) x standard price
Shampoo = {(310*1.3) - 705} x $0.30 = $91 Unfavorable

Water = (310*11 - 6100)*0.04 = $108 Unfavorable

2. Labor rate variance = (Standard rate - actual rate) x Actual hours used
= {$9 - ($2600/250)} x 250 = $350 Unfavorable

Labor efficiency variance = (Standard hours - actual hours) x Standard rate
= {(310*0.85) - 250} x $9 = $122 Favorable